### v1

@@ 3.2 Extrinsic sensitivity
where is the source’s production time (when activity was measured by a detector by the Radiology department staff of the Hôpitaux Univérsitaires de Genève (HUG), with a precision down to tens of kBq), is the half-life of metastable Tc-99, is the area of the FOV, and is the surface of the Flood Phantom. The sensitivity over time is therefore computed as:
The reported error on the sensitivity measurements is defined as the quadrature sum of the statistical uncertainty on the rate and the systematic uncertainty on the activity, giving the uncertainty in sensitivity. The latter is the propagated error of the initial injected activity in the phantom (set at of the initial activity) and the uncertainty on the custom-made phantom’s dimension:
The procedure is repeated with the LEHS collimator and the LEHR collimator, placing the source at 2cm from the camera’s lid.
3.3 Energy resolution
@@
3.4 Collimator septal penetration
The camera’s external shielding was tested to assess the leakage of gamma rays from the Tungsten collimator’s walls, which are designed to prevent external radiation from reaching the scintillator crystal. An IGC does not require strong directional sensitivity, unlike 1D gamma probes, which rely on it due to their lack of imaging capability. Directionality is, therefore, a key performance aspect typically studied in detail for probes [Wydra2005]. In the case of POSiCS, a leakage study was performed to ensure that possible activity hotspots outside the FOV do not influence the imaging performance of the POSiCS camera.
This test was performed by registering the detected rate of a Co-57 point source in free air, positioned 3 cm from the center of the camera, at three distinct angles: one at relative to the normal of the FOV, and the other two at and from the same normal direction, as shown in Figure 4. The 100-140 keV energy window (used throughout this work for Co-57 data processing) was adopted.
Rate data was collected by registering 150 packets of 4000 events for each under-test angle. For each packet, the initial and final timestamps are used to compute the average rate of that packet. The 150 rates per angle are presented in Figure 9, normalized by the mean value of the rate at from the FOV normal direction.
3.5 Energy response to other radiotracers (Lu-177)
@@ 4.1 Extrinsic spatial resolution
By placing the source in contact with the camera, we obtained a resolution with the LEHR collimator of 1.4 0.1 mm, which increased to 2.8 0.1 mm when the source was moved 3 cm away. In contact with the camera head, the LEHR and LEHS collimators show comparable resolutions. However, due to the well-known collimator projective effect [cherry_ch14, AngerCamera], the slope of the resolution over distance for the LEHS collimator shows a much steeper trend. With the latter collimator, we measured a spatial resolution of 1.9 0.1 mm at 0 cm from the source, which increased to 5.1 0.1 mm at 3 cm. Moreover, the difference between measurements in free air and with scatterer material is within error bars using the LEHR collimator, while the presence of PMMA worsens resolution with the LEHS collimator. This phenomenon is likely related to the collimator’s improved ability to reject photons scattered in the PMMA layers between the source and the camera.
@@ 4.2 Extrinsic sensitivity
| Sensitivity results | LEHS collimator | LEHR collimator |
@@
For the LEHS mode camera, we report a sensitivity of after rejecting all events whose energy does not lie within the spectral window adopted in this work for Tc-99m (120 keV-160 keV). Instead, we observe a significant reduction in sensitivity for the LEHR collimator, down to . We identify the reason for the mismatch between theoretical predictions and the actual detected sensitivity (about 4%-20%, depending on the collimator) as being related to the choice of energy window containing photoelectric absorption events and the uncertainty the image correction algorithm adds.
The results of the sensitivity evaluation are displayed in Figure 7 for both collimators. For the camera in HR mode, we observed a measured sensitivity distributed as a Gaussian function around its central value, with a standard deviation of 6 cps/MBq. This is displayed in the sensitivity projection on the right side-plot. The Gaussianity of this distribution implies that the negative exponential decay correction to the activity matches the observed rate. Furthermore, from Table 2, we evince that the Skewness and Kurtosis of the points distribution (side plot) are compatible with a Gaussian distribution, for which the typical skewness is 0 and the typical kurtosis is 3. For the LEHS collimator, we observe a distribution of sensitivities characterized by a slight positive skewness (indicating a heavier right tail) and a kurtosis slightly below 3, which suggests a more peaked distribution. This is also evident in the time series of measured sensitivities, which shows a positive vertical bending of the estimated sensitivity around 600 s (cf. Figure 7). We identify this problem with the high count rate registered with the high-sensitivity collimator, suggesting that saturation effects are beginning to build up. Therefore, the results we obtained should be considered a lower limit on the sensitivity, which could be higher if a lower-activity source had been placed in front of the camera.
@@ 4.3 Energy resolution
The achieved energy resolution with the LEHS collimator is 19.14%, which aligns with the typical energy resolutions reported in the literature for SFOV gamma cameras [Farnworth2023]. However, this result is relatively sharp for a LYSO-based device [Nakanishi2017, Deprez2014, Morozov2017]. This energy resolution is enough to distinguish the photopeak from the nearby structures in the energy spectrum, namely the backscattering peak (located at about 90 keV, as expected) and the Compton edge: the third peak from the Right in Figure 8.
With the LEHR collimator, it was possible to measure an energy resolution of 19.94%, which, as expected, is comparable with the resolution achieved through the LEHS collimator. Energy resolution is primarily driven by the scintillator (the number of photons produced per scintillation event), the optical coupling to the photodetector, and the photodetector efficiency. Therefore, as assessed by these tests, the collimator does not significantly impact the energy resolution. However, we notice a large amount of scattered gamma rays with the LEHR collimator, which we link to the increased amount of high-Z scatterer material (tungsten) present due to the longer collimator.
4.4 Collimator septal penetration
The tests with both collimators show that the count rate drops significantly when the source is placed outside of the FOV. Moreover, the response is symmetric when placing the source outside of the FOV to the left and Right of the camera. In the case of the LEHS collimator, the average rate drop outside the FOV is 88%, indicating that the shielding around the scintillator effectively eliminates most of the incoming gamma rays. The LEHR collimator shows an average relative rate drop of only 61%. Since both collimators have the same shielding thickness, the observed discrepancy may suggest that the greater amount of tungsten in the LEHR collimator leads to increased gamma-ray scattering. As a result, Compton-scattered gamma rays may still reach the scintillator and be detected. This implies that the energy window cut applied around the photopeak is less effective when using the LEHR collimator.
@@ 4.5 Camera’s response to other radiotracers (Lu-177)
The obtained energy spectrum is displayed in Figure 10. The main emission peak was found to be at 218 keV, which shows, therefore, a 4.6% discrepancy with the theoretical value (208 keV). This behaviour is expected, since the camera was calibrated based on a 140 keV reference energy, and the non-linearity of the camera’s light detector charge response creates a slight shift in the central energy value. For this reason, the energy look-up table (LUT) was rescaled to ensure that the main emission peak was centered at 208 keV. With this correction, it was possible to identify the peaks presented in Table 3.
@@
We recognize the first peak as very low-energy X-rays and background radiation. We identify peak number 2 as the X-ray emissions at 55 keV and 64 keV, which merge into a single peak at 60 keV due to the limited energy resolution of the camera at such low energies, as already observed by [Sagisaka2024]. Peak number 3 is the 113 keV emission, which is clearly distinguishable and well centered around its reference value. We identify peak number 4 as the Compton shoulder of the main emission line (peak 5). It is also possible to notice a small bump in the spectrum after the 208 keV energy, which could correspond to a weak emission of Lu-177 (321 keV). However, this energy has a low photoelectric absorption probability within the thin scintillator crystal of the POSiCS camera and, therefore, cannot be identified. Moreover, this higher-energy region is also populated by the Lu-176 background due to the presence of this isotope in the LYSO:Ce scintillator (cf. subsection 2.3).
| Peak number | Central energy | Energy resolution |
| Peak 1 | 34 keV | - |
| Peak 2 | 60 keV | - |
| Peak 3 | 113 keV | 35% |
| Peak 4 | 163 keV | - |
| Peak 5 | 208 keV | 19% |
@@
The primary emission energies used for Lu-177 SPECT imaging are the 113 keV and 208 keV gamma-rays. In our case, the energy resolution of these two photopeaks was determined through a Gaussian fit and found to be 35% and 19%, respectively. We therefore proceeded with the usual calibration pipeline, by defining two 20% energy windows around these peaks (following the approach of [Marin2017]) and a large energy window (0-76 keV) to assess the reconstructed images through low-energy X-ray emission.
@@
The imaging results after the standard image correction pipeline, Gaussian kernel convolution (following the NEMA standards guidelines [NEMA2023]), and spline interpolation, are presented in Figure 11. The cavity image can be efficiently reconstructed for both the X-ray and 113 keV energy windows. However, we encounter significant problems in reliably reconstructing an image at 208 keV. As the Figure 11 shows, gamma-ray detection in the 208 keV range extends beyond the cavity’s region of interest (ROI), with signals detected in pixels outside the area filled with activity. This issue arises from the use of thin collimator septa, which are optimized for lower energies and fail to constrain non-perpendicular gamma rays at 208 keV adequately.
@@
This test demonstrated that the POSiCS camera, without any modification to its configuration or components, is also suitable for imaging with Lu-177. It effectively detects events associated with the isotope’s X-ray emissions as well as gamma emissions up to its 133 keV peak. However, imaging at higher energies would require the use of a longer collimator. The performance of the current LEHR collimator with Lu-177 is presently under investigation.
4.6 Performance parameters discussion
@@
| Spatial resolution AEC (122 keV) @0cm [mm] | 1.9 0.1 | 1.4 0.3 |
| Spatial resolution AEC (122 keV) @3cm [mm] | 5.1 0.1 | 2.8 0.3 |
@@
Table 4 provides a summary of the key performance parameters of the POSiCS camera. The results confirm that the system fulfills its intended purpose, enabling imaging with millimeter-scale spatial resolution: a critical feature for intraoperative applications. Notably, when equipped with the LEHS collimator, the camera achieves high sensitivity compared to other intraoperative gamma systems used for radioguided surgery (RGS) [Farnworth2023]. This is particularly advantageous for rapid image acquisition, minimizing disruptions to the surgical workflow.
@@
The measured energy resolution is generally broader compared to other scintillator cameras with higher light yields, such as LaBr3:Ce or the more commonly used NaI:Tl cameras (or if compared to direct-detection gamma cameras, such as CZT-based devices). Nevertheless, the observed resolution remains competitive, especially when compared to larger LYSO-based systems operating near 140 keV, which often exhibit energy resolutions exceeding 20%. Moreover, LYSO offers additional benefits: its high density allows for the development of compact detectors with enhanced sensitivity.
Shielding tests revealed some residual gamma-ray penetration from the sides and indicated significant scattering associated with the LEHR collimator. This is also evidenced by the corresponding energy spectrum acquired with the same scintillator (see Figure 8). An improved collimator design with enhanced lateral shielding could further mitigate this scattering and leakage.
5 Outlook and future developments

### v2

@@ 2.2.2 Sensitivity
where is the source’s production time (when activity was measured by a detector available at Geneva University Hospital, with a precision down to tens of kBq), is the half-life of metastable Tc-99m, is the area of the FOV, and is the surface of the Flood Phantom. The sensitivity over time is therefore computed as:
Here, is the trigger rate of the camera at time . The reported error on the sensitivity measurements is defined as the quadrature sum of the statistical uncertainty on the rate and the systematic uncertainty on the activity, giving the uncertainty in sensitivity. The latter is the propagated error of the initial injected activity in the phantom (set at of the initial activity) and the uncertainty on the custom-made phantom’s dimension:
The procedure is repeated with the LEHS collimator and the LEHR collimator, placing the source at 2cm from the camera’s lid. No background subtraction was performed, due to the high activity of the source with respect to the 500 cps rate from LYSO:Ce, which is reduced to few tens of cps after energy window selection.
2.2.3 Energy resolution
@@
2.2.4 Collimator side penetration
The camera’s external shielding was tested to assess the penetration of gamma rays from the Tungsten collimator’s walls and the septal penetration, which are designed to prevent external radiation from reaching the scintillator crystal. An IGC does not require strong directional sensitivity, unlike 1D gamma probes, which rely on it due to their lack of imaging capability. Directionality is, therefore, a key performance aspect typically studied in detail for probes [Wydra2005]. In the case of POSiCS, a penetration study was performed to ensure that possible activity hotspots outside the FOV do not influence the imaging performance of the POSiCS camera. Moreover, penetration of gamma-ray from the side are ultimately visually identified and thus can be removed by truncating the image.
This test was performed by registering the detected rate of a Co-57 point source in free air, positioned 3 cm from the center of the camera, at three distinct angles: one at relative to the normal of the FOV, and the other two at and from the same normal direction. The 100-140 keV energy window (used throughout this work for Co-57 data processing) was adopted.
The rate data was collected in 150 packets of 4000 events for each tested angle. For each packet, the initial and final timestamps are used to compute the average rate of that packet.
2.2.5 Energy response to other radiotracers (Lu-177)
@@ 3.1 Extrinsic spatial resolution
By placing the source in contact with the camera, we obtained a resolution with the LEHR collimator of 1.4 0.1 mm, which increased to 2.8 0.1 mm when the source was moved 3 cm away. In contact with the camera head, the LEHR and LEHS collimators show comparable resolutions. However, due to the well-known collimator projective effect [cherry_ch14, AngerCamera], the resolution of the LEHS collimator degrades faster than for the LEHR collimator. With the LEHR collimator, we measured a spatial resolution of 1.9 0.1 mm at 0 cm from the source, which increased to 5.1 0.1 mm at 3 cm. Moreover, the difference between measurements in free air and with scatterer material is within error bars using the LEHR collimator, while the presence of PMMA worsens resolution with the LEHS collimator. This phenomenon is likely related to the collimator’s improved ability to reject photons scattered in the PMMA layers between the source and the camera.
@@ 3.2 Extrinsic sensitivity
| Sensitivity results | LEHS coll. | LEHR coll. |
@@
For the LEHS mode camera, we report a sensitivity of after rejecting all events whose energy does not lie within the spectral window adopted in this work for Tc-99m (120 keV-160 keV). Instead, we observe a significant reduction in sensitivity for the LEHR collimator, down to , as expected.
The results of the sensitivity evaluation are displayed in Figure 6 for both collimators. For the camera in HR mode, we observed a measured sensitivity distributed as a Gaussian function around its central value, with a standard deviation of 8 cps/MBq. The mean and the standard deviation were weighted by the points variance. This is displayed in the sensitivity projection on the right side-plot. The Gaussianity of this distribution implies that the negative exponential decay correction to the activity matches the observed rate. This suggests that the fluctuations after the source’s activity compensation are mostly driven by poissonian shot-noise. Furthermore, from Table 2, we evince that the Skewness and Kurtosis of the points distribution (side plot) are compatible with a Gaussian distribution, for which the typical skewness is 0 and the typical kurtosis is 3. For the LEHS collimator, we observe a distribution of sensitivities characterized by a slight positive skewness (indicating a heavier right tail) and a kurtosis slightly below 3, which suggests a more peaked distribution. This is also evident in the time series of measured sensitivities, which shows a positive vertical bending of the estimated sensitivity around 600 s (cf. Figure 6). Since the kurtosis and skewness analyses still suggest a mostly Gaussian behaviour, we cannot explain this visual bending at this stage as a systematic feature. On the other hand, it is possible that saturation effects due to the high count rate introduce systematic distortions in the time series. This feature will be further investigated. Due to the possibility of saturation, especially at the beginning of the acquisition time (when the source activity is maximal), we consider the measured extrinsic resolution of the LEHS collimator as a lower limit.
@@ 3.3 Energy resolution
The achieved energy resolution with the LEHS collimator is 19.14% 0.26%, which aligns with the typical energy resolutions reported in the literature for SFOV gamma cameras [Farnworth2023]. However, this result is relatively sharp for a LYSO-based device [Nakanishi2017, Deprez2014, Morozov2017]. This energy resolution is enough to distinguish the photopeak from the nearby structures in the energy spectrum, namely the backscattering peak (located at about 90 keV, as expected) and the Compton edge: the third peak from the Right in Figure 7.
With the LEHR collimator, it was possible to measure an energy resolution of 19.94% 0.20%, which, as expected, is comparable with the resolution achieved through the LEHS collimator. Energy resolution is primarily driven by the scintillator (the number of photons produced per scintillation event), the optical coupling to the photodetector, and the photodetector efficiency. Therefore, as assessed by these tests, the collimator does not significantly impact the energy resolution. However, we notice a large amount of scattered gamma rays with the LEHR collimator, which we link to the increased amount of high-Z scatterer material (tungsten) present due to the longer collimator.
3.4 Collimator lateral leakeage
The tests with both collimators show that the count rate drops significantly when the source is placed outside of the FOV. Moreover, the response is symmetric when placing the source outside of the FOV to the left and Right of the camera. In the case of the LEHS collimator, the average rate drop outside the FOV is 88%, indicating that the shielding around the scintillator effectively eliminates most of the incoming gamma rays. The LEHR collimator shows an average relative rate drop of only 61%. Since both collimators have the same shielding thickness, the observed discrepancy may suggest that the greater amount of tungsten in the LEHR collimator leads to increased gamma-ray scattering. As a result, Compton-scattered gamma rays may still reach the scintillator and be detected. This implies that the energy window cut applied around the photopeak is less effective when using the LEHR collimator. The rate as function of angle is shown in Figure 8, relative to the the mean rate at from the FOV normal direction.
@@ 3.5 Camera’s response to other radiotracers (Lu-177)
The obtained energy spectrum is displayed in Figure 10. The main emission peak was found to be at 218 keV, which shows, therefore, a 4.6% discrepancy with the theoretical value (208 keV). This behaviour is expected, since the camera was calibrated based on a 140.5 keV reference energy, and assumes linear response of the light yield to the reconstructed charge. For this reason, the energy look-up table (LUT) was rescaled to ensure that the main emission peak was centered at 208 keV.
@@
We recognize the first peak as very low-energy X-rays and background radiation, coming from electronics noise and characteristic X-rays from the lead shielding. We identify peak number 2 as the X-ray emissions at 55 keV and 64 keV, which merge into a single peak at 60 keV due to the limited energy resolution of the camera at such low energies, as already observed by [Sagisaka2024]. Peak number 3 is the 113 keV emission, which is clearly distinguishable and well centered around its reference value. We identify peak number 4 as the Compton shoulder of the main emission line (peak 5). It is also possible to notice a small bump in the spectrum after the 208 keV energy, which could correspond to a weak emission of Lu-177 (321 keV). However, this energy has a low photoelectric absorption probability within the thin scintillator crystal of the POSiCS camera and, therefore, cannot be identified. Moreover, this higher-energy region is also populated by the Lu-176 background due to the presence of this isotope in the LYSO:Ce scintillator (cf. subsection 2.1.2).
@@
The primary emission energies used for Lu-177 SPECT imaging are the 113 keV and 208 keV gamma-rays. In our case, the energy resolution of these two photopeaks was determined through a Gaussian fit and found to be 35% and 19%, respectively. We therefore proceeded with the usual reconstruction pipeline, by defining two 20% energy windows around these peaks (following the approach of [Marin2017]) and a large energy window (0-76 keV) to assess the reconstructed images through low-energy X-ray emission. Those peaks were already taken as a reference for imagin with Lu-177 in other works [Roth2020, Roth2024].
@@
The imaging results after the standard image correction pipeline, Gaussian kernel convolution (following the NEMA standards guidelines [NEMA2023]), and spline interpolation, are presented in Figure 9 (more information about the correction pipeline are reported in subsection 2.1). The cavity image can be efficiently reconstructed for both the X-ray and 113 keV energy windows. However, we encountered significant problems in reliably reconstructing an image at 208 keV. As shown in Figure 9, gamma-ray detection in the 208 keV range extends beyond the cavity’s region of interest (ROI), with signals detected in pixels outside the area filled with activity. This issue arises from the use of thin collimator septa, which are optimized for lower energies and fail to collimate gamma rays at 208 keV adequately.
@@
This test demonstrated that the POSiCS camera, without any modification to its configuration or components, is also suitable for imaging with Lu-177. It effectively detects events associated with the isotope’s X-ray emissions as well as gamma emissions up to its 113 keV peak. However, imaging at higher energies would require the use of a high energy collimator with longer septa. The performance of the current LEHR collimator with Lu-177 is presently under investigation.
4 Discussion
4.1 Comparison with existing literature
@@
| Spatial resolution AEC (122 keV) @0cm [mm]∗ | 1.9 0.1 | 1.4 0.3 |
| Spatial resolution AEC (122 keV) @2cm [mm]∗ | 4.2 0.1 | 2.4 0.3 |
@@
In this work, we presented a performance study of the newly developped POSiCS camera. A summary table of these results is presented in Table 3.
To reference POSiCS performances within the landscape of already existing IGCs, a comparison with the literature is proposed, starting with the following parameters: Spatial resolution, extrinsic sensitivity, energy resolution and device’s weight.
The evaluation of POSiCS performance parameters can be compared with literature using the comprehensive 2023 review on IGCs [Farnworth2023]. From this review, the relevant performance and physical characteristics of other intraoperative SFOV gamma cameras were extracted. The mean value, the standard deviation, and the minimal and maximal values of these parameters are reported in Table 4. Only devices for which the target parameter was available were included in the statistical analysis and comparison. For cameras featuring multiple operating modes (e.g., interchangeable collimators of different lengths), each parameter was selected from the configuration that most favored the corresponding performance metric, to be conservative. The devices parameters distribution compared to the results of this work for the POSiCS camera are presented in Figure 11 and commented below.
For spatial resolution, only devices for which the resolution at 0 cm from the source was specified were taken into consideration, in order to be compared with the in-contact performances of POSiCS. It is important to highlight that the evaluation of POSiCS spatial resolution was performed at an energy of 122 keV, while some of the devices used for the comparison evaluated the same parameter at 140 keV. POSiCS shows an extrinsic resolution of a few millimeters even at distances of 4 cm from the target source. This requirement is satisfied with both collimators. The resolution of the camera at 0 cm from the source outperforms 9 devices out of 10 in both LEHR and LEHS modes. However, this trend may vary at longer source-to-camera distances due to the geometrical projection effect of the collimator. Overall, POSiCS meets the millimetric spatial resolution requirement up to a source-to-collimator distance of 4 cm, in agreement with the initial specifications for SLNB and respecting the requirements for SLNB.
Since the extrinsic sensitivity of a gamma camera is less dependent on the distance from the source [cherry_ch14], we opted to consider all the devices with known sensitivity estimation, regardles of the reference source-to-collimator distance from the source at which the parameter was estimated.
From Table 4, we notice a large standard deviation of the devices’ sensitivity statistics, which is driven by the different purposes of the cameras presented in the 2023 IGC review (some of which target high resolution, and others high sensitivity). The LEHR camera outperforms only 2 out of 15 devices in terms of sensitivity, since the main target of this configuration is to have a high-precision in image reconstruction, and is therefore more suited for higher activity sources. The LEHS mode, instead, shows a sensitvity well above average, outperforming 12 devices out of 15. Only one device has a significantly higher sensitivtiy (about 1500 cps/MBq). This specific gamma camera leverages a peculiar collimator-scintillator configuration enhancing the device’s sensitivity, and represents a far outlier in the statistics.
@@
The measured energy resolution of the POSiCS camera is in line with the average of 21.4% presented in the review (even if this value, once again, was evaluated at 122 keV and 140 keV, depending on the device). The cameras showing a significantly sharper energy resolution with respect to POSiCS are either scintillator cameras based on scintillators with higher light yields, such as LaBr3:Ce or IGCs performing imaging with a direct detection system, such as CZT detectors that are therefore more suited for spectroscopy. Nevertheless, the observed energy resolution remains competitive, especially when compared to larger LYSO-based systems operating near 140 keV, which often exhibit energy resolutions exceeding 20%. Moreover, LYSO offers additional benefits: its high density allows for the development of compact detectors without reducing the devices’ sensitivity.
A comparison of the cameras’ weights was also performed. This parameter is of great importance to introduce gamma cameras within the protocol of SNLB and ROLL standard procedures, since it allows a easier manipulation of the device. POSiCS outperforms 9 devices out of 10 with the LEHR collimator configuration, while it outperforms all the devices in LEHS mode, with a weight well below the average of 1.27 kg. Only one device had a weight comparable to POSiCS (about 320 g), but with a FOV of 1.31.3 cm2, implying a very small device and hence a reduced weight.
4.2 Lateral collimator leakage tests
Shielding tests revealed some residual gamma-ray penetration in both collimator. When outside of the FOV, the camera mounted with the LEHS collimator showed a rate drop of 88%, while in LEHR configuration, the rate drop was of only 61%. We explain this due to the increased amount of scattering associated to the longer collimator (as proven by the full FOV energy spectrum of the camera), and to the lower amount of accepted photons by the LEHR collimator. The violin plots in Figure 8 show that the distribution of normalized rates is broader for the HR collimator than for the HS configuration. This larger variability is observed because the HR collimator registers a lower absolute rate, as it is designed to reject a larger fraction of incoming trajectories (i.e. reject more gamma-rays). Since the leakage from the camera’s sides is expected to be similar for both collimators (given their identical shielding thickness) the same absolute leakage produces a larger relative contribution in the HR case. As a result, the relative rate reduction is smaller for the HR collimator. This implies that the lateral leakage has a stronger impact on image reconstruction for the LEHR configuration, and might dominate when observing low activity sources within the FOV (for instance when observing near the injection zone).
The analysis shows that an improved collimator design with enhanced lateral shielding could mitigate this penetration effects, and that might be required especially for the HR collimator.
4.3 POSiCS performances at different energies
Tests with 177Lu were performed due to the increased need for radiation monitoring in 177Lu-DOTA-TATE and 177Lu-PSMA procedures, and to test POSiCS performances at different energies.
The tests revealed that it was possible to image a liquid 177Lu source in a plastic mouse phantom, with the two energy windows: 0-76 keV and 90-136 keV, targeting low energy Lu-177 emissions (X-Rays) with the first window and the 113 keV emission line, with the second window. The third emission line, at 208 keV, while composing the larger fraction of detected gamma-rays, was not suited for imaging. The collimator walls of the LEHS collimator were too thin to correctly absorb rejected trajectory, and this results in a strong leakage in the whole camera FOV, as shown in Figure 9.
Tests with the LEHR collimator are currently being performed and the possibility of a Lu-177 based collimator with thicker septa is under evaluation.
The higher septal penetration due to the 208 keV emission of Lu-177 is a well known problem in SFOV gamma cameras, and has been reported for other devices [Roth2020, Roth2024]. Once again, the assestment and evaluation of a Lu-177 dedicated collimator could allow the POSiCS camera to provide imaging at 208 keV with the simple need of swapping the device’s head.
| Parameter | # Devs | Mean | Std | Min | Max |
| Spatial resolution [mm] | 10 | 2.44 | 0.95 | 1.0 | 4.5 |
| Sensitivity [cps/MBq] | 15 | 317 | 356 | 21 | 1500 |
| Energy resolution [%] | 15 | 21.3 | 13.7 | 5.2 | 58 |
| Weight [kg] | 10 | 1.27 | 0.76 | 0.32 | 3.2 |
5 Conclusions